# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

### v5

@@ Abstract
The origin of life is often approached through the lens of replication, heredity, or molecular specificity. This paper proposes a thermodynamic framework in which the emergence of life is driven by the persistence of reaction pathways that align energetically with fluctuating environmental inputs. We define a reaction viability inequality based on energy input, return, resilience, and expenditure, which selects for persistent chemical configurations without invoking heredity or genetic encoding.
We further incorporate entropic dynamics and spatial constraints into an augmented persistence function, showing that systems far from equilibrium can simultaneously increase global entropy while supporting localized chemical order. These refinements lead to the development of the Thermodynamic Abiogenesis Likelihood Model (TALM), a probabilistic extension that estimates the likelihood of persistence-driven selection under various prebiotic and planetary scenarios.
This framework redefines the conditions under which life-like organization may emerge and provides a testable, general theory for abiogenesis grounded in physical law.
@@ 1 Introduction
The origin of life remains one of the most profound and unresolved questions in science. Traditional approaches emphasize replication, heredity, and biochemical specificity under non-equilibrium thermodynamic conditions as prerequisites for natural selection. The transition from abiotic chemistry to biological evolution is often framed in terms of autocatalytic reaction networks sustained by external energy sources. However, the mechanism by which such systems become selectively retained remains underdeveloped in many existing models.
@@
Here, we propose a general theoretical framework in which selection-like behavior can emerge in the absence of heredity or replication–driven instead by thermodynamic constraints on energy input, use, storage, and dissipation. This model formalizes a statistical pathway to abiogenesis rooted in the persistence of reaction networks whose viability depends on environmental energy availability, internal energy dynamics, and structural resilience. By reframing natural selection as a consequence of energy-compatible persistence, we unify entropy production, spatial organization, and reaction viability under a physically grounded, molecule-agnostic theory applicable across planetary environments. In this formulation, natural selection is not exclusive to genetics but can emerge in any system where energy regulation enables sustained reaction cycles.
@@
This work contributes a scalable thermodynamic model that reconceptualizes the origin of life as the emergence of persistent chemical structure within energy-modulated environments. It requires no initial heredity, replication, or carbon-based chemistry. Grounded in first principles yet agnostic to molecular identity, the theory offers a generalizable framework for how abiogenesis may arise from prebiotic reaction chains shaped by entropy flux and diffusion constraints.
@@
We extend this framework into a probabilistic formulation, the Thermodynamic Abiogenesis Likelihood Model (TALM), which estimates the likelihood of persistence-driven selection under diverse planetary conditions. TALM captures the emergence of selective behavior in localized chemical environments and scales across planetary and cosmic contexts. The remainder of this paper develops the model’s thermodynamic foundation, introduces its statistical structure, and explores its implications across both Earth-like and unconventional life-permitting scenarios.
@@
As a test of the model’s predictive scope, we propose a minimal comparative system involving amphiphilic molecules of varying chain length. This system isolates entropy and structural persistence as primary variables and allows for experimental validation of TALM’s core mechanisms under well-defined aqueous conditions.
@@ 2 Terminology and Variable Definitions
To aid clarity and interdisciplinary accessibility, the following table summarizes key variables and terms used in the model:
Symbol / Term Description
$k$ The number of reactions in the reaction chain
$z(t)$ Time-varying environmental input energy available to the system. Reflects fluctuating external sources such as thermal, photonic, or chemical energy.
$y(t)$ Stored energy available within the system. Acts as a buffer during energy deficits and enhances persistence through time-lagged supply.
$x_{i}$ The energy required to perform reaction $i$.
$r_{i}$ The energy released during reaction $i$.
$s_{i}$ The energy stored in reaction products.
$\sum r_{i}$ Total energy released as usable output from individual reactions within the system. Represents internal efficiency.
$\sum x_{i}$ Total energy required to maintain all reactions in the system. Represents the system’s energetic load or metabolic cost.
$R_{n}$ Structural or network-based resilience term. Reflects robustness of the reaction topology, such as redundancy, feedback loops, or autocatalytic features.
$P_{E}$ Probability that the environment provides survivable energy fluctuations.
$P_{C}$ Probability that sufficient chemical complexity or molecular diversity exists.
$P_{P}$ Probability that reaction chains can persist over time through internal energy logic.
$P_{S}$ Probability that the system achieves dynamic stability after initial fluctuations.
$P_{life}$ Composite probability that all four preconditions are satisfied for life-like persistence.
TALM Thermodynamic Abiogenesis Likelihood Model. A generalized framework to assess the likelihood of persistent prebiotic systems under dynamic energetic constraints.
Persistence The ability of a system to sustain itself over time, despite energy fluctuations, without requiring replication or heredity.
Table 1: Summary of key variables and terms in the TALM framework.
3 Core Thermodynamic Framework
Thermodynamic Assumptions and Principles
@@ 3 Core Thermodynamic Framework
This model is constructed on a set of thermodynamic principles and assumptions about how energy is used, transferred, and stored in prebiotic chemical systems. These assumptions are not intended to capture the full complexity of biochemistry, but rather to provide a minimalist framework capable of illustrating how selection-like dynamics may arise from energy-driven processes alone.
Energy Availability and Use
@@
Dynamic Environmental Input
We introduce a time-dependent energy function, $z(t)$, representing the usable energy made available from the environment. This can fluctuate due to factors such as sunlight cycles, geothermal activity, or chemical gradients. This treatment reflects empirical work showing that surface thermal cycling, particularly in hydrothermal field environments, can generate temporally structured energy availability conducive to prebiotic polymerization (Deamer, 2011) (Ross D & Deamer D, 2019). The persistence of a reaction chain is therefore not only a function of its internal energy efficiency but also its ability to maintain viability under changing energy input.
Reaction Chains as Systems
@@
Stored Energy and Future Use
Energy stored in products ($s_{i}$) is not immediately usable unless another reaction releases it. Thus, stored energy acts as a potential buffer or reserve, contributing to a system’s future viability but not necessarily to its current energy budget.
System Definition
Let the system consist of a series of $n$ reactions, each requiring a defined amount of energy to proceed and potentially releasing or storing energy in the process. We define the following terms:
  • $k$: The number of reactions in the chain
  • $x_{i}$: The energy required to perform reaction $i$.
  • $r_{i}$: The energy released during reaction $i$.
  • $s_{i}$: The energy stored in reaction products.
  • $y(t)$: Residual energy at time $t$.
The Reaction Viability Inequality

### v7

# A Thermodynamic Model for the Emergence of Natural Selection in Prebiotic Reaction Networks

@@ Abstract
The origin of life is often approached through the lens of replication, heredity, or molecular specificity. This paper proposes a thermodynamic framework in which the emergence of life is driven by the persistence of reaction pathways that align energetically with fluctuating environmental inputs. We define a reaction viability inequality based on energy input, release, resilience, and expenditure, which selects for persistent chemical configurations without invoking heredity or genetic encoding. We further incorporate entropic dynamics and spatial constraints into an augmented persistence function, showing that systems far from equilibrium can simultaneously increase global entropy while supporting localized chemical order. These refinements lead to the development of the Thermodynamic Abiogenesis Likelihood Model (TALM), a probabilistic extension that estimates the likelihood of persistence-driven selection under diverse prebiotic and planetary scenarios. This framework redefines the conditions under which life-like organization may emerge and provides a testable, general theory for abiogenesis grounded in physical law.
keywords:
Abiogenesis; Thermodynamic selection; Prebiotic chemistry; Entropy; Reaction networks; Origin of life
@@ 1 Introduction
The origin of life remains one of the most challenging problems in science. While considerable progress has been made in identifying plausible prebiotic chemistries and environmental settings, a general physical principle that explains how lifelike organization can first emerge remains elusive. Existing models variously emphasize autocatalytic networks [1], dissipative adaptation [2], kinetic stability [3], and amphiphilic self-assembly into membranes and compartments [4, 5]. Each of these approaches captures an aspect of the problem, but none provide an operational criterion for assessing the persistence of chemical systems in fluctuating prebiotic environments.
Recent theoretical work [6] has examined the thermodynamic limits of Darwinian evolution in molecular replicators, deriving bounds that relate energy dissipation to differential reproductive success. While that analysis addresses systems in which heredity and competition are already established, the present study concerns a distinct and antecedent regime, one in which persistence itself functions as a selection filter prior to the emergence of replication. In this sense, the two approaches are complementary. Kolchinsky’s work formalizes thermodynamic constraints on post-replicative evolution, while the present model examines how selection-like [^1] In this work the term selection-like refers solely to differential persistence arising from thermodynamic compatibility in fluctuating environments and does not imply adaptive fitness, heredity, or biological evolution. dynamics could arise spontaneously from energy-structure interactions in prebiotic chemistry.
@@
Here we propose a theoretical framework in which selection-like behavior emerges without heredity or replication, driven instead by thermodynamic constraints on energy input, use, storage, and dissipation. This model formalizes a statistical pathway to abiogenesis rooted in the persistence of reaction networks whose viability depends on environmental energy availability, internal energy dynamics, and structural resilience. This formulation treats natural selection as a consequence of energy-compatible persistence. It links entropy production, spatial organization, and reaction viability within a physically grounded, molecule-agnostic framework. In this view, selection-like behavior can emerge in any system where energy regulation supports sustained reaction intervals.
@@
This work develops a thermodynamic model in which the origin of life is understood as the emergence of persistent chemical structure within energy-modulated environments. The model assumes no initial heredity, replication, templating, enzymatic catalysis, or specific molecular identity. Grounded in first principles and agnostic to molecular identity, the framework describes how abiogenesis may arise from prebiotic reaction chains shaped by entropy flux and diffusion constraints.
@@
We extend this framework into a probabilistic formulation, the Thermodynamic Abiogenesis Likelihood Model (TALM), which estimates the likelihood of persistence-driven selection under diverse planetary conditions. TALM characterizes the emergence of persistence-based selection in localized chemical environments and can be extended to diverse planetary settings. The remainder of this paper develops the model’s thermodynamic foundation, introduces its statistical structure, and examines its implications for prebiotic chemical organization.
@@
As an initial test of the model’s predictions, we analyze a comparative system involving amphiphilic molecules of varying chain length. This system isolates entropy and structural persistence as primary variables and allows for experimental validation of the core mechanisms under well-defined aqueous conditions.
@@ 2 Terminology and Variable Definitions
To aid clarity and interdisciplinary accessibility, the following table summarizes key variables and terms used in the model [^2] Molecular-scale values are expressed in J per molecule; values in the main text use J·mol–1 for clarity, convertible via Avogadro’s number.:
Table 1: Symbols, units, and descriptions used in the framework.
Symbol Units Description
$k$ dimensionless Number of reactions in the reaction chain.
$z(t)$ J·mol-1 Time-varying environmental energy input available to the system.
$y(t)$ J·mol-1 Residual energy at time $t$; positive values indicate persistence.
$x_{i}$ J·mol-1 Energy required to perform reaction $i$ (activation or maintenance cost).
$r_{i}$ J·mol-1 Energy released as usable output from reaction $i$.
$s_{i}$ J·mol-1 Potential stored energy in reaction products that can contribute to $S(t)$.
$S(t)$ J·mol-1 Stored energy within the system; evolves according to $\frac{dS}{dt}=\sum_{i}s_{i}-U(S,t)$.
$U(S,t)$ J·mol-1·s-1 Release or utilisation rate of stored energy as a function of $S(t)$ and time.
$R_{n}$ J·mol-1 Resilience contribution, defined as $R_{n}=\lambda_{R}\mathcal{R}(G)$, where $\mathcal{R}(G)\!\in\![0,1]$ is a normalized redundancy factor from network topology and $\lambda_{R}$ is an energy scale (typically $RT$ per mole).
$\Delta S_{\text{net}}$ J·K-1·mol-1 Net entropy change ($\Delta S_{\text{env}}-\Delta S_{\text{local}}$) between system and environment.
$T\Delta S_{\text{net}}$ J·mol-1 Entropy term expressed in energy units at temperature $T$.
$D(x,t)$ m2·s-1 Diffusion coefficient describing spatial dispersion or confinement.
$\alpha$ J·mol-1·K Scaling coefficient for the entropy-related penalty contribution.
$\beta$ J·mol-1·s·m-2 Scaling coefficient for the diffusion-related penalty contribution.
$\Phi(t)$ J·mol-1 Entropic-diffusive penalty term $\alpha_{S}T\Delta S_{\text{local}}(t)+\beta_{D}\dfrac{RT}{D_{0}}D(x,t)$.
$y^{\prime}(t)$ J·mol-1 Augmented persistence function $y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$.
$P_{\text{selective}}$ dimensionless Probability of persistence across $N$ reaction permutations: $P_{\text{selective}}=1-(1-P(y^{\prime}(t)\geq 0))^{N}$.
[…]
3 The Abiogenesis Persistence Framework
3.1 Assumptions and Principles
@@
This model is constructed on a set of thermodynamic principles and underlying assumptions about how energy is used, transferred, and stored in prebiotic chemical systems. These assumptions are not intended to capture the full complexity of biochemistry, but rather to provide a minimalist framework capable of illustrating how selection-like dynamics may arise from energy-driven processes alone.
3.2 Defining Persistence
Persistence, as used in this model, is not introduced as a fundamental thermodynamic state variable. Instead, it represents an emergent property of a driven chemical system: the ability to maintain non-zero reaction fluxes, concentration gradients, or structural organisation over time in the presence of dynamic environmental input [7]. Such persistence reflects the system’s sustained capacity to export entropy to its surroundings and to maintain a favourable balance between energy uptake, dissipation, and loss [8]. Related concepts such as dynamic kinetic stability similarly treat extended molecular or network lifetimes as consequences of energy flux and reaction kinetics rather than as primitive thermodynamic variables [3]. Experimental studies on wet–dry cycling and protocell formation likewise show that structural persistence in prebiotic contexts depends critically on continuous environmental driving [9]. In this framework, persistence serves as a practical, derived descriptor for assessing whether prebiotic reaction systems remain viable over time.
3.3 Energy Availability and Use
@@
3.4 Dynamic Environmental Input
We introduce a time-dependent energy function, $z(t)$, representing the usable energy made available from the environment. This can fluctuate due to factors such as sunlight cycles, geothermal activity, or chemical gradients. This treatment reflects empirical work showing that surface thermal cycling, particularly in hydrothermal field environments, can generate temporally structured energy availability conducive to prebiotic polymerization [5] [10]. The persistence of a reaction chain is therefore not only a function of its internal energy efficiency but also its ability to maintain viability under changing energy input.
3.5 System Definition
Let the system consist of a series of $k$ reactions, each requiring a defined amount of energy to proceed and potentially releasing or storing energy in the process. We define the following terms:
  • $k$: The number of reactions in the chain
  • $x_{i}$: The energy required to perform reaction $i$.
  • $r_{i}$: The energy released during reaction $i$.
  • $s_{i}$: The potential stored energy in reaction products.
  • $S(t)$: Stored free energy within the system.
  • $y(t)$: Residual energy at time $t$.
3.5.1 Reaction Chains as Systems
@@
3.5.2 The Reaction Viability Inequality